# Can Transformers Do Enumerative Geometry?

### v1

@@ 5. Methods
Dynamic Range Activator. It has been shown [parascandolo2017taming, dauphin2017language, ziyin2020neural, so2021searching] that the choice of activation functions plays an important role in the interpolation and extrapolation properties of Transformers and, more generally, neural networks. To enable Transformer models to precisely capture the recursive behavior of the data, we introduce a simple activation function that we call the Dynamic Range Activator (DRA). DRA integrates both harmonic and hyperbolic components, as follows,
@@
where $a,b,c,d$ are learnable parameters. This diversity allows the function to simultaneously model periodic data (through sine and cosine), rapid growth or attenuation (through the hyperbolic tangent), and maintain a linear baseline response. DRA is inspired by the Snake non-linear function [ziyin2020neural]. However, Snake only offers a sinusoidal modulation added to a linear term, which, while providing a basic non-linear transformation, lacks the additional flexibility for rapid damping or amplification effects that hyperbolic tangents can provide. To demonstrate the advantage of DRA in capturing recursive behavior, we set up a small experiment. We generate a small dataset based on the recursive function $r(n)=n+(n,\mathrm{AND},r(n-1))$, where AND is the bitwise logical AND operator [10.1007]. We train a fully connected neural network with one hidden layer consisting of $64$ neurons over the interval $n\in[0,150]$, then test the model over $n\in[151,200]$. The results are shown in figure 3.
Figure 3. Comparison of DRA with Snake, ReLU and Tanh non-linear activations in training region $n\in[0,150]$ and test (extrapolative) region $n\in[151,200]$.
@@
This experiment demonstrates that neither ReLU nor Tanh is capable of capturing the recursive nature of the underlying function, as both baselines perform naive interpolation and extrapolation. Snake appears to learn some periodicity in both the training and testing regions. In contrast, DRA shows a better ability to capture the correct fluctuations, as well as the rapid growth and drops of the underlying recursive function, in both the interpolation and extrapolation regimes. This provides evidence that DRA has the desired flexibility for modeling recursive behavior and has the potential to effectively model such problems. We further conduct experiments to demonstrate the advantages of DRA over typical non-linear activation functions in predicting $\psi$-class intersection numbers.
[DYN] registry tokens and Talking Modalities. In multi-modal learning, effectively sharing and integrating information across different modalities is crucial for building robust models. Inspired by the success of Self-Supervised Learning (SSL) [gui2024survey], we introduce a simple modification to enhance correspondence between modalities. To achieve this, we concatenate Dynamic tokens, [DYN], to each modality. These tokens are similar to [CLS] tokens in BERT [devlin2019bert] and Vision Transformers [dosovitskiy2021], in that they are passed through the attention layers alongside the input data. However, unlike [CLS] tokens, they are not directly used for the downstream supervised task of amplitude prediction. Instead, they are used indirectly, like register tokens [darcet2024vision], which store and process global information in Vision Transformers.
By attending to all tokens in the input, [DYN] tokens, similar to register tokens, facilitate a soft contrastive learning update, enabling them to learn a global context for each sample. Unlike register tokens, [DYN] tokens are also optimized to ensure that the outputs from different modalities of data—specifically from the $B$ tensor, which provides global information per sample, and the partitions $\bm{d}$, which serve as local information carriers—are as similar as possible while simultaneously reducing the redundancy of the information they carry. Their optimization is guided by the Canonical Correlation Analysis [Abdi2018] family of SSL, where the aim is to infer the relationship between two modalities by analyzing their cross-covariance matrices [balestriero2023cookbook]. Inspired by Barlow Twins [zbontar2021barlow], the two modalities start to have a “conversation”, promoting an interaction in which the two modalities collaboratively refine their representations. The idea behind this is that such a consensus mechanism not only enhances the model’s ability to generalize better to Out-Of-Distribution (OOD) data but also improves its robustness. The Talking Modalities loss is then defined as follows:
$\mathcal{L}_{\textup{TM}}=\sum_{i=1}^{k}\left(1-\mathcal{C}_{ii}^{(\bm{B},\bm{d})}\right)^{2}+\lambda\sum_{\begin{subarray}{c}i,j=1,\ldots,k\\ i\neq j\end{subarray}}\left(\mathcal{C}_{ij}^{(\bm{B},\bm{d})}\right)^{2},$ (5.2)
where $\mathcal{C}$ is the cross-correlation matrix of the representations from the two modalities, $(\bm{B},\bm{d})$, $k$ is the dimensionality of the model’s embeddings (vector representations), and $\lambda$ is the weighting factor that balances the contributions of talking modalities. The first term penalizes the diagonal elements of the cross-correlation matrix $\mathcal{C}$ for deviating from the identity, thus encouraging invariance. The second term penalizes the off-diagonal elements, encouraging the components of the representation vectors to be decorrelated. In section 6, we will show that this modification enhances the performance of the model in the OOD prediction regime.
After the Transformer trunk of the network, all the embedded information is aggregated by a permutation-invariant block. We employ the Principal Neighbourhood Aggregation (PNA) layer [corso2020principal] to pool information from the modalities $(\bm{B},\bm{d})$. Once the information from these two branches, for each sample, is modulated with their corresponding genus and number of marked points positional attributes, $(g,n)$, it is mapped to the outputs, i.e. $\psi$-class intersection numbers, via an MLP head. DRA is the non-linear activation function used throughout these blocks. As the loss function, we use the Mean Absolute Error (MAE) loss. Given that the range of amplitudes for the intersection numbers is very wide, $[10^{-45},10^{45}]$, MAE was the most appropriate metric. Essentially, MAE is not disproportionately affected by the massive range of the target values, unlike Mean Squared Error (MSE), which can be overly influenced by the squared terms of large numbers. Thus, the total loss is as follows:
$\mathcal{L}_{\textup{Total}}=\mathcal{L}_{\textup{MAE}}+\mathcal{L}_{\textup{TM}}\,.$ (5.3)
6. Results
To evaluate the performance of our approach, we consider two setups: In-Distribution (ID) and Out-Of-Distribution (OOD) settings. In these settings, we face both length extrapolation and context extrapolation challenges. Given the wide range of values for $\psi$-class intersection numbers, we use R-squared ($\mathrm{R}^{2}$) to compare the results and assess the goodness-of-fit in the intersection number regression task. $\mathrm{R}^{2}$ reflects the proportion of the variance in the dependent variable that is predictable from the independent variables. The higher the value of $\mathrm{R}^{2}$, the greater the relationship between prediction and ground truth values. For instance, an $\mathrm{R}^{2}$ value of $0.75$ indicates that 75% of the total variation in the data is explained by the model, while the remaining 25% is not captured by the model.
Uncertainty Quantification. To estimate the uncertainty of the prediction, we incorporate Conformal Prediction (CP) [vovk1999machine]. CP is a probabilistic uncertainty quantification technique that provides prediction intervals in finite samples without making any distributional assumptions. We integrate Inductive Conformal Prediction [4410411] with a dynamic sliding window [xu2023conformalpredictiontimeseries] that computes rolling residuals for predicted intersections for each set of marked points $n$. In our modification, we fit a quantile regression model to the residuals within a rolling window of marked points, allowing us to predict the desired quantile that adapts to local variations in the residuals. Given the high heteroscedasticity in the data, where the variance of residuals changes across intersections, we employ a sliding window over the interval of each set of marked points. This ensures that the residuals are more scale-dependent and appropriately adjusted for the inherent variability in the data, potentially capturing trends and heteroscedasticity better. As a result, we report the Normalized Conformity Score (NCS)
$\mathrm{NCS}\coloneqq\frac{|\langle\bm{d}\rangle_{g,n}-\langle\bm{d}\rangle_{g,n}^{\textup{pred}}|}{\langle\bm{d}\rangle_{g,n}}$ (6.1)
and the Conformal Interval Width (CIW).
We perform two sets of experiments. First, we report the performance of DynamicFormer in both the ID and OOD settings. We compare DRA with Snake [ziyin2020neural] and GLU [so2021searching] activation functions as baselines to validate DRA’s non-linearity. We also compare the results with and without the [DYN] register token [darcet2024vision] and the Talking Modalities mechanism in the OOD setting. Thereafter, we engage in an interesting discussion on interoperability, correlational, and causal analysis of what the model is learning. We also explore how to perform abductive hypothesis testing based on the model’s internal understanding of the underlying enumerative problem to prove conjectures.
6.1. In-Distribution Results
@@
In the ID setting, we examine data with the same genera as the training data, that is $g_{\textup{ID}}=14$, but with different, unseen numbers of marked points $n_{\textup{ID}}\in[3,9]$. In this setting, the $\mathrm{R}^{2}$ and NCS are shown in table 1. It is noteworthy that this is a test dataset where the model has not seen any of these amplitudes. The comparison between the ground truth data and the predicted values is shown in figure 4.
The average CIW for the ID setting is $0.03$, with an empirical coverage of $79\%$. This empirical coverage falls short of our target coverage of $90\%$, indicating that our conformal procedure is under-covering. This shortfall could be due to the severe recursive nature of the data, suggesting that our calibration data is still too localized and therefore not fully representative of the entire test domain. Despite this limitation, the $79\%$ coverage is the highest we have achieved with our conformal calibration method. In comparison, other methods, such as naive Inductive Conformal Prediction [4410411], achieved only $53\%$ coverage, and Conformalized Quantile Regression [romano2019conformal] yielded an empirical coverage of $57\%$.
Table 1. $\mathrm{R}^{2}$ and NCS results with DRA activation function in the ID setting.
$(g,n)$ $\mathrm{R}^{2}\uparrow$ NCS $\downarrow$
$(14,3)$ $97.2\pm 0.1$ 0.19
$(14,4)$ $98.4\pm 0.1$ 0.11
$(14,5)$ $99.3\pm 0.7$ 0.08
$(14,6)$ $98.1\pm 0.5$ 0.13
$(14,7)$ $99.0\pm 0.8$ 0.09
$(14,8)$ $99.9\pm 0.3$ 0.01
$(14,9)$ $99.9\pm 0.3$ 0.01
Figure 4. Comparison of true, predicted, and $\mathrm{R}^{2}$ values in the ID setting. The plot shows the true values, the predicted values, and the $\mathrm{R}^{2}$ scatter plot, including a line representing perfect prediction.
6.2. Out-Of-Distribution Results
@@
Now, we study the more challenging task of OOD prediction. In the OOD setting, we examine data with higher genus than the training data, specifically $g_{\textup{OOD}}=[15,16]$, and a number of marked points $n_{\textup{OOD}}\in[2,7]$. In this setting, the $\mathrm{R}^{2}$ metric is shown in table 2. The average CIW for the OOD setting is $1.45\times 10^{-5}$, with an empirical coverage of $76\%$. The comparison between the ground truth data and the predicted values is shown in figure 5.
The Snake and other activation functions failed completely in this OOD setting, as shown in table 3. These results demonstrate that DynamicFormer can capture intersection numbers even for genera and numbers of marked points outside the training data. Table 4 compares the performance of the model in the OOD regime with and without the [DYN] register token (i.e. the multi-modal version of [REG]) and the Talking Modalities mechanism. This experiment shows that the [DYN] register token generally provides a small boost to performance, and adding the Talking Modalities mechanism further improves prediction accuracy. We did not observe any performance improvement over the training data or in the ID setting. It seems that using register tokens as a method to query and process global information helps with the model’s generalization abilities.
Table 2. $\mathrm{R}^{2}$ and NCS results with DRA activation function in the ODD setting.
$(g,n)$ $\mathrm{R}^{2}\uparrow$ NCS $\downarrow$
$(15,2)$ $82.6\pm 0.9$ 0.27
$(15,3)$ $97.7\pm 0.8$ 0.11
$(15,4)$ $98.4\pm 0.9$ 0.11
$(15,5)$ $99.4\pm 0.2$ 0.08
$(15,6)$ $98.4\pm 0.8$ 0.13
$(15,7)$ $98.5\pm 0.3$ 0.11
$(16,3)$ $98.1\pm 0.6$ 0.13
$(16,4)$ $80.0\pm 0.4$ 0.44
$(16,5)$ $98.9\pm 0.3$ 0.10
Figure 5. Comparison of true, predicted, and $\mathrm{R}^{2}$ values in the ODD setting. The plot shows the true values, the predicted values, and the $\mathrm{R}^{2}$ scatter plot, including a line representing perfect prediction.
As evidenced by table 3, we see that the DRA outperforms the GLU and Snake activation functions by a large margin in predicting the recursive intersection numbers in the OOD setting, where the models have to regress intersection numbers with previously unseen numbers of marked points.
@@
Table 3. Comparison of $\mathrm{R}^{2}$ and CIW between Transformer-based models with GLU, Snake, and DRA as their activation functions in the OOD regime $g\in[15,16]$ and $n\in[2,7]$.
GLU Snake DRA
$\mathrm{R}^{2}\uparrow$ $\text{CIW}^{(\times 10^{-5})}$downarrow $\mathrm{R}^{2}\uparrow$ $\text{CIW}^{(\times 10^{-5})}$downarrow $\mathrm{R}^{2}\uparrow$ $\text{CIW}^{(\times 10^{-5})}$downarrow
$84.1\pm 0.4$ 9.33 $68.9\pm 0.8$ 200 $\mathbf{95.8}\pm 0.6$ $\mathbf{1.45}$
Table 4. Comparison of $\mathrm{R}^{2}$ and CIW between models with and without [DYN] register token and Talking Modalities. A small performance improvement can be observed as a result of using register tokens Talking Modalities in the OOD regime $g\in[15,16]$ and $n\in[2,7]$.
None [DYN] [DYN]+TM
$\mathrm{R}^{2}\uparrow$ $\text{CIW}^{(\times 10^{-5})}$downarrow $\mathrm{R}^{2}\uparrow$ $\text{CIW}^{(\times 10^{-5})}$downarrow $\mathrm{R}^{2}\uparrow$ $\text{CIW}^{(\times 10^{-5})}$downarrow
$91.1\pm 0.4$ $1.95$ $94.6\pm 0.6$ $1.46$ $\mathbf{95.8}\pm 0.6$ $\mathbf{1.45}$
6.3. Insights, Interoperability and Discussion
@@
Up until now, we have tried to showcase that DynamicFormer, as a modified Transformer, is capable of performing enumerative geometry and predicting $\psi$-class intersection numbers. However, several key questions remain. Do Transformers actually understand the underlying mathematics? How does the AI agent learn this computation? Is it possible to extract useful knowledge or hints toward a possible (re)discovery of a theorem?
To address these questions, we have made several interesting observations throughout our work that offer deeper insights into the model’s behavior and learning process, hinting at potential mathematical knowledge discovery.
6.3.1. The Dilaton Equation
@@
The topological recursion formula (3.8) for the specific case of $\psi$-class intersection numbers and the choice of $d_{1}=1$ takes a particularly simple form, known as the Dilaton equation. After relabelling, it reads
@@
Let $\langle\bm{d}\rangle_{g,n}$ be the intersection numbers indexed by genus $g$ and number of marked points $n$. Given the the set of values of the tensor $\bm{B}$ and the set of partitions $\bm{d}\in\mathbb{N}^{n}$, the neural network embedding $p_{g,n}\colon\set{\mathbb{R}^{d_{g,n}\times d_{g,n}\times d_{g,n}},\mathbb{N}^{n}}\rightarrow\mathbb{R}^{k}$ is a vector space of dimension $k$ and is denoted as $p_{g,n}(\langle\bm{d}\rangle_{g,n}\mid\bm{B},\bm{d})$. This embedding describes the hidden understanding of the Transformer (right before the prediction head). If we normalize these embeddings to unit length and then compute the self-projection of these embedding vectors onto themselves, we obtain the cosine of the angle between the normalized embeddings of $\psi$-class intersection numbers as a measure of similarity:
$S_{i,j}=\frac{\langle p_{g,n}^{i},p_{g,n}^{j}\rangle}{\lVert p_{g,n}^{i}\rVert\,\lVert p_{g,n}^{j}\rVert}\,.$ (6.3)
The similarity matrix $S_{i,j}$ represents the geometric proximity between two pairs of $\psi$-class intersection numbers. Values in this matrix range from $-1$ to $1$, where $-1$ denotes complete dissimilarity, $1$ indicates perfect similarity, and 00 suggests orthogonality or decorrelation.
Figure 6. The Dilaton relations matrix (left) between $\psi$-classes intersection numbers for $g=14$ (top) and $g=15$ (bottom). Cosine similarity between embeddings of predicted intersection numbers (right). Cosine similarity between embeddings of predicted intersection numbers with a cut $S_{i,j}\geq 0.98$ (middle).
Looking at the cosine similarity between samples in the embedding space of the model’s internal activations, shown in figure 6, we observed some interesting recursive patterns. We hypothesize that these patterns stem from the Virasoro constraints between intersection numbers. To verify this, we visualized the Dilaton equations and confirmed that the model has effectively rediscovered it as relationships between intersection numbers. We translated the Dilaton equation into a Dilaton relation matrix by flagging any two intersection numbers that satisfy this equation. We then observed that the intersection numbers with high cosine similarity follow at least the Dilaton equation. We anticipate that other observed patterns may also be manifestations of the remaining Virasoro constraints. This phenomenon persists even in the OOD setting where the model has not seen any samples. This may explain how the model has succeeded in generalizing well by learning the underlying symmetries and relations. This provides evidence that DynamicFormer has independently learned the the Virasoro constraints in a data-driven way, without prior exposure to these governing rules.
6.3.2. Large genus limit and abductive reasoning
Delecroix–Goujard–Zograf–Zorich [delecroix2021masur] conjectured that, in the large genus limit, $\psi$-class intersection numbers simplify to a specific asymptotic form. This conjecture was subsequently proven by Aggarwal [aggarwal_large], incorporating a combinatorial and probabilistic analysis of the associated Virasoro constraints.

### v3

@@ 3 Methods
Dynamic Range Activator. It has been shown (Parascandolo et al., 2017; Dauphin et al., 2017; Ziyin et al., 2020; So et al., 2021) that the choice of activation functions plays an important role in the interpolation and extrapolation properties of Transformers. To enable Transformer models to precisely capture the recursive behavior of the data, we introduce a simple activation function that we call the Dynamic Range Activator (DRA). DRA integrates both harmonic and hyperbolic components as follows,
@@
where $a,b,c,d$ are learnable parameters. It allows the function to simultaneously model periodic data (through sine and cosine) and rapid growth or attenuation (through the hyperbolic tangent) response. DRA is inspired by the Snake non-linear function (Ziyin et al., 2020). However, Snake only offers a sinusoidal modulation added to a linear term, which, while providing a basic non-linear transformation, lacks the additional flexibility for rapid damping or amplification effects that hyperbolic tangents can provide.
@@
To demonstrate the advantage of DRA in capturing recursive behavior, we set up a small experiment. We generate a small dataset based on the recursive function $r(n)=n+(n,\mathrm{AND},r(n-1))$, where AND is the bitwise logical AND operator (Sloane, 2007). We train a fully connected neural network with two hidden layers consisting of $64$ and $32$ neurons over the interval $n\in[0,120]$, then test the model over $n\in[121,200]$. we also provide a comparison between Multi-Layer Perceptron (MLP) and the vanilla Kolmogorov–Arnold Networks (KAN) (Liu et al., 2024). The results are shown in Figure 1. This experiment demonstrates that ReLU, Tanh, Gated Linear Unit (GLU) (So et al., 2021), and KAN are unable to capture the recursive nature of the underlying function within a finite training time. Snake appears to learn some periodicity in both the training and testing regions. In contrast, DRA shows a better ability to capture the correct fluctuations, as well as the rapid growth and drops of the underlying recursive function, in both the interpolation and extrapolation regimes. This provides an evidence that DRA has the desired flexibility for modeling recursive behavior and has the potential to effectively model such problems. We further conduct experiments, in Section 4, to demonstrate the advantages of DRA over typical non-linear activation functions in predicting $\psi$-class intersection numbers.
Figure 1: Comparison of DRA MLP with Snake, ReLU, Tanh, GLU, Swish, and PolyNorm activation functions, and KAN in training region $n\in[0,120]$ and test (extrapolative) region $n\in[121,200]$. We ensured that all models have a comparable number of parameters. The MLP with DRA non-linearity demonstrates superior performance in capturing the function’s behavior.
4 Results
To evaluate the performance, we consider two setups: In-Distribution (ID) and Out-Of-Distribution (OOD) settings. We use $\mathrm{R}^{2}$ to compare the results and assess the goodness-of-fit in the intersection number regression task. We then compare DRA with ReLU, Snake, and GLU activation functions as baselines.
4.1 In-Distribution Results
@@
In the ID setting, we examine data with the same genera as the training data, that is $g_{\textup{ID}}=[1,13]$, but with different, unseen numbers of marked points $n_{\textup{ID}}\in[35,11]$. In this setting, the $\mathrm{R}^{2}$, the empirical coverage, and the conformal width (CW) (see Appendix G) are shown in Table 1(a).
$(g,n)$ $\mathrm{R}^{2}\uparrow$ Coverage CW
$(1,35)$ $99.8$ 90.35 1.03
$(2,33)$ $99.6$ 83.60 0.84
$(3,31)$ $99.9$ 74.79 0.76
$(4,29)$ $98.7$ 95.66 1.11
$(5,27)$ $99.1$ 92.66 1.03
$(6,25)$ $99.3$ 91.18 0.80
$(7,23)$ $99.1$ 93.05 0.68
$(8,21)$ $99.8$ 90.88 0.76
$(9,19)$ $99.9$ 96.71 0.91
$(10,17)$ $99.9$ 90.01 1.04
$(11,15)$ $99.8$ 91.97 0.87
$(12,13)$ $99.6$ 89.08 1.30
$(13,11)$ $99.9$ 95.90 0.97
(a) $\mathrm{R}^{2}$ and conformal uncertainty estimation results with $\alpha=0.1$ ($90\%$ target coverage) in the ID setting.
$(g,n)$ $\mathrm{R}^{2}\uparrow$ Coverage CW
$(14,[1,9])$ $99.6$ 93.82 0.93
$(15,[1,7])$ $95.9$ 84.27 0.91
$(16,[1,5])$ $94.1$ 89.60 3.55
$(17,[1,3])$ $93.8$ 95.27 8.30
(b) $\mathrm{R}^{2}$ and conformal uncertainty estimation results with $\alpha=0.1$ ($90\%$ target coverage) in the OOD setting.
Table 1: $\mathrm{R}^{2}$ and conformal uncertainty estimation results.
4.2 Out-Of-Distribution Results
@@
Now, we study the more challenging task of OOD prediction. In the OOD setting, we examine data with a higher genera than the training data, specifically $g_{\textup{OOD}}=[14,15,16,17]$, and a number of marked points $n_{\textup{OOD}}\in[1,9]$. In this setting, the $\mathrm{R}^{2}$, the empirical coverage, and the conformal width are shown in Table 1(b). As evidenced by Table 2, the DRA outperforms the ReLU, GLU, and Snake activation functions in predicting the recursive intersection numbers in the OOD setting.
@@
Table 2: Comparison of $\mathrm{R}^{2}$ and Conformal Width between models with ReLU, GLU, Snake, and DRA as their activation functions in the OOD regime.
ReLU GLU Snake DRA
$\mathrm{R}^{2}\uparrow$ CW $\downarrow$ $\mathrm{R}^{2}\uparrow$ CW $\downarrow$ $\mathrm{R}^{2}\uparrow$ CW $\downarrow$ $\mathrm{R}^{2}\uparrow$ CW $\downarrow$
$71.5$ 9.73 $74.7$ 8.34 $82.9$ 6.55 $\mathbf{95.8}$ $\mathbf{3.42}$
5 How Does the Network Do Enumerative Geometry?
@@
Up until now, we have tried to showcase that DynamicFormer is capable of predicting $\psi$-class intersection numbers. However, several key questions remain. Do Transformers actually understand the underlying enumerative geometry? Is it possible to extract useful knowledge or hints toward a possible (re)discovery of a theorem? To address these questions, we have made several interesting observations throughout our work that offer deeper insights into the network’s “world model” and learning process, hinting at potential mathematical knowledge discovery.
5.1 The Dilaton Equation
@@
The topological formula equation 2.4 for the specific case of $\psi$-class intersection numbers and the choice of $d_{1}=1$ takes a particularly simple form, known as the Dilaton equation,
@@
Let $\langle\bm{d}\rangle_{g,n}$ be the intersection numbers indexed by genus $g$ and number of marked points $n$. Given the the set of values of the tensor $\bm{B}$ and the set of partitions $\bm{d}\in\mathbb{N}^{n}$, the neural network embedding $p_{g,n}\colon\set{\mathbb{R}^{d_{g,n}\times d_{g,n}\times d_{g,n}},\mathbb{N}^{n}}\rightarrow\mathbb{R}^{k}$ maps the input data to a $k$ dimensional vector space $\mathbb{R}^{k}$ denoted as $\bm{x}_{g,n}(\langle\bm{d}\rangle_{g,n}\mid\bm{B},\bm{d})$. This embedding describes the hidden understanding of the Transformer just before the prediction head, capturing the learned representation of the intersection numbers. By equipping $\mathbb{R}^{k}$ with the standard inner product, $S_{i,j}:\mathbb{R}^{k}\times\mathbb{R}^{k}\rightarrow\mathbb{R}$, we obtain the cosine of the angle between the normalized embeddings of $\psi$-class intersection numbers as a measure of geometric similarity:
$S_{i,j}=\frac{\langle x_{g,n}^{i},x_{g,n}^{j}\rangle}{\lVert x_{g,n}^{i}\rVert\,\lVert x_{g,n}^{j}\rVert}\,.$ (5.2)
Observing an interesting recursive pattern in Figure 2, we hypothesize that they stem from the Virasoro constraints between intersection numbers. To verify this, we visualized the Dilaton equations and confirmed that the model has rediscovered it as relationships between intersection numbers. We translated the Dilaton equation into a Dilaton relation matrix by flagging any two intersection numbers that satisfy this equation. We then observed that the intersection numbers with high cosine similarity follow at least the Dilaton equation. We anticipate that other observed patterns may also be manifestations of the remaining Virasoro constraints. This phenomenon persists even in the OOD setting. This explains how the model has succeeded in generalizing well by learning the underlying symmetries and relations. It provides an evidence that the network has learned the Virasoro constraints in a data-driven way without prior exposure to these governing rules.
Figure 2: The Dilaton relations matrix (left) for $\psi$-class intersection numbers at $g=13$(top) and $g=14$(bottom). Cosine similarity between embeddings of predicted intersection numbers (right) and with a cut $S_{i,j}\geq 0.98$ (middle).
5.1.1 Large Genus Asymptotic and Abductive Reasoning
Delecroix et al. (2021) conjectured that, in the large genus limit, $\psi$-class intersection numbers simplify to a specific asymptotic form.